I need some help with this problem.
The speed of a mototboat in still water is 15 mi/hr. If it takes the motorboat the same time to go 15 mi upstream the river as it takes to go 30 mi
downstream, what is the speed of the current?
Thank You!

Answer by stanbon(75887)

The speed of a mototboat in still water is 15 mi/hr. If it takes the motorboat the same time to go 15 mi upstream the river as it takes to go 30 mi
downstream, what is the speed of the current?
--------------
Let the rate of the current be "x":
Upstream DATA:
Distance = 15 mi ; rate = 15-x ; time = d/r = 15(15-x) mph
-----------
Downstream DATA:
Distance = 30 mi ; rate = 15+x ; time = d/r = 30/(15+x) mph
-------------
EQUATION:
time up = time down
15/(15-x) = 30/(15+x)
Divide both sides by 15 to get:
1/(15-x) = 2/(15+x)
15+x = 2(15-x)
15+x = 30-2x
3x=15
x=5 mph
The speed of the current is 5 mph
---------------
Cheers,
Stan H.